# AttentionStore: Cost-effective Attention Reuse across Multi-turn Conversations in Large Language Model Serving

### v2

@@ 6 Conclusion
This paper proposes AttentionStore, a new attention that allows the reuse of the KV caches for any ensuing turns of the same conversation, achieving a significant reduction in the recomputation overhead of KV caches in LLMs. To improve the efficiency of AttentionStore, we design overlapped KV cache access, hierarchical KV cache placement, and positional encoding decoupled KV cache truncation schemes. Extensive experimental results demonstrate that AttentionStore significantly decreases the TTFT by up to 87% and improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations. It reduces the end-to-end inference cost by up to 70%. It also decreases the TTFT by up to 95% and enhances prompt prefilling throughput by 22$\times$ for long sequence inference.

### v3

@@ 6 Conclusion
This paper proposes CachedAttention, a new attention that reuses the KV caches for any ensuing turns of the same conversation, achieving a significant reduction in the recomputation overhead of KV caches in LLMs. To improve the efficiency of CachedAttention, we design overlapped KV cache access, hierarchical KV cache placement, and positional encoding decoupled KV cache truncation schemes. Extensive experimental results demonstrate that CachedAttention significantly decreases the TTFT by up to 87% and improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations. It reduces the end-to-end inference cost by up to 70%.
Acknowledgments
This work was partly supported by Huawei Cloud, the Advanced Research and Technology Innovation Centre (ARTIC) at the National University of Singapore under grant FCT-RP1 A-0008129-00-00, and the Ministry of Education in Singapore grants A-0008143-00-00 and A-0008024-00-00.